OpenStudy (anonymous):

solve the equation in the set of complex numbers. give only exact answers, not approximations... 4x(x-2)-3x(x+1)=-18

OpenStudy (anonymous):

x=2,9

OpenStudy (anonymous):

x=1,10

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

4x^2-8x-3x^2-3x+18=0 x^2-11x+18=0 (x-2)(x-9)=0 so, x=2,9

OpenStudy (anonymous):

dont know how luckey got different answer..

OpenStudy (anonymous):

yeah thats why i needed the work haha

OpenStudy (anonymous):

yeah 2,9 only

OpenStudy (anonymous):

that was a calculation mistake